Consider the following numerical model:

C = 700 + 0.8YD

I = 500-2000i+0.1Y

G = 400

T = 500

- Find the equation for the IS curve (keep in mind that YD=Y-T). If you were to draw the IS curve, what would the intercept (where the IS curve crosses the Y vertical axis) be? (and find the intercept in relation to Y - ie in the form Y=a-bi where a is the intercept)

- If G increases from 400 to 600, how much income (Y) changes for any given level of the interest rate in the goods-market equilibrium?

- If T decreases from 500 to 300, how much income (Y) changes for any given level of the interest rate in the goods-market equilibrium? (Add a + or - to the number you reach depending on whether the change is positive or negative)
Answer #1

10 +07 (-) ya Despesabls T500 & yd 6-500) nome >> C= 700 +0.8LY-SPD) 333333333 15: ly=C+I+G (Goods maxkat equilib Y 700 +0.8LAY = 0.8 AY TO AY + AG (from equation (D) ised only those terme in which AG is useful AY= 0.8 Ay+oAY + AG AY-0.90Y DG OLAY =